OpenStudy (anonymous):

Can I get the formula needed to solve this problem? Find the slope of a line parallel to the line through the given points. A (3, 8), B(-2, 5)

OpenStudy (helder_edwin):

the formula of the slope is \[ \large \frac{y_0-y_1}{x_0-x_1} \]

OpenStudy (amistre64):

to avoid mixing up the values i just stack the points and subtract

OpenStudy (amistre64):

A (3, 8) - B(-2, 5) ---------- 5 , 3 ; stack y/x for the slope ... 3/5

OpenStudy (amistre64):

use that in the slope point formula to define the eqation with \[y-y_0=m(x-x_0)\] where m is the slope and (x0,y0) is either point

OpenStudy (anonymous):

Would it be 0.6? Or 3/5

OpenStudy (amistre64):

those are the same value arent they? and for some reason i read the question as finding an equation ... lol

OpenStudy (amistre64):

lets say 2 feet = 24inches is it right to say something measures 2 feet? or, 24 inches?

OpenStudy (anonymous):

Well, I'm simply wondering if the answer was correct, to make sure I understand how the formula works.

OpenStudy (amistre64):

oh, then yes; it can be written as either .6 or 3/5 then ... depends on whos grading it maybe

OpenStudy (anonymous):

Alright, thanks guys!